OpenStudy (anonymous):

Find the perimeter and area of a right triangle if one leg measures 27 yd. and the other leg measures 36 yd. Students must show all work step by step to receive full credit for the problem. Include correct units with your solution.

OpenStudy (anonymous):

x^2+y^2=z^2 27^2+36^2=z^2 z=45 P=x+y+z P=27+36+45 P=108 is this correct

OpenStudy (anonymous):

Yes, it is.